# GeoPT: Scaling Physics Simulation via Lifted Geometric Pre-Training

### v1

@@ Appendix G Full Results
To supplement the scaling results in Fig. 6 of the main text, we further test the corresponding performance in DTCHull and Car-Crash in Table G. It is also observed that pre-training with GeoPT can help avoid potential overfitting, especially in industrial design tasks where only limited data is available. Besides, GeoPT can also take advantage of diverse geometry and dynamics conditions, highlighting the value of rich 3D geometry assets.
@@
Here, we present the concrete values in Tables 6-10 for five simulation tasks in Fig. 1 and 5. We color some results to highlight GeoPT’s capability in improving performance, convergence and reducing data requirements:
@@
Due to space limitations in the main text, we present the complete visualization of the physics states in Fig. 21-24 as a supplement to Fig. 3 and 7, which includes the learned correlations under various supervisions and dynamics-dependent correlations prompted by different configurations of dynamics conditions.
@@
Appendix H Limitations and Future Work
@@
GeoPT primarily targets simulations involving complex geometries, a common requirement in industrial design and a setting where neural simulators have achieved their most notable successes. Additionally, it is also interesting to see the extension of GeoPT to simulations without complex geometry boundaries, such as 3D turbulence in regular grids (Perlman et al., 2007). Since such simulations are mainly about computationally intensive physics interactions, the previous explorations (Holzschuh et al., 2025) still rely on expensive physics supervision. Taking such regular grid simulation into account, one possible way is to leave a certain ratio of pre-training iterations for dynamics under an empty geometry boundary during the pre-training process of GeoPT, where all the tracking points are in free-flight motion, which can make the model learn the intrinsic isotropic dynamics. We would like to leave this exploration as future work.

### v2

@@ Appendix F Full Results
To supplement the scaling results in Fig. 6 of the main text, we further test the corresponding performance in DTCHull and Car-Crash in Table F. It is also observed that pre-training with GeoPT can help avoid potential overfitting, especially in industrial design tasks where only limited data is available. Besides, GeoPT can also take advantage of diverse geometry and dynamics conditions, highlighting the value of rich 3D geometry assets.
@@
Statistical test
In Table 7, we also include the standard deviation of key results under 5 runs for statistical analysis. All training-data-saving claims are supported by 95% confidence intervals, where the equivalence margin is set as 0.001.
Table 7: Standard deviation of relative L2 under 5 runs.
Standard deviation ($\times 10^{-3}$) DrivAerML NASA-CRM AirCraft DTCHull Car-Crash
From Scratch 100% Training data 0.43 0.94 1.03 1.34 0.71
GeoPT 100% Training data 0.68 0.35 1.23 0.88 0.31
GeoPT 80% Training data 0.80 0.66 1.27 1.10 0.56
GeoPT 60% Training data 0.81 0.93 1.17 0.49 0.13
GeoPT 40% Training data 0.31 0.76 1.83 0.91 0.76
@@
Here, we present the concrete values in Tables 8-12 for five simulation tasks in Fig. 1 and 5. We color some results to highlight GeoPT’s capability in improving performance, convergence and reducing data requirements:
@@
Due to space limitations in the main text, we present the complete visualization of the physics states in Fig. 22-25 as a supplement to Fig. 3 and 7, which includes the learned correlations under various supervisions and dynamics-dependent correlations prompted by different configurations of dynamics conditions.
@@
Appendix G Limitations and Future Work
@@
GeoPT primarily targets simulations involving complex geometries, a common requirement in industrial design and a setting where neural simulators have achieved their most notable successes. Additionally, it is also interesting to see the extension of GeoPT to simulations without complex geometry boundaries, such as 3D turbulence in regular grids (Perlman et al., 2007). Since such simulations are mainly about computationally intensive physics interactions, the previous explorations (Holzschuh et al., 2026) still rely on expensive physics supervision. Taking such regular grid simulation into account, one possible way is to leave a certain ratio of pre-training iterations for dynamics under an empty geometry boundary during the pre-training process of GeoPT, where all the tracking points are in free-flight motion, which can make the model learn the intrinsic isotropic dynamics. We would like to leave this exploration as future work.